OpenStudy (anonymous):

What is 4x + 2y = -20 written in slope-intercept form?

OpenStudy (anonymous):

y = 2x - 20 y = -4x + 20 y = 20x - 4 y = -2x - 10

OpenStudy (anonymous):

solve for y: \[4x + 2y = -20\] \[2y=-20-4x\] \[y=-10-2x\] or \[y=-2x-10\]